[figure: 9.7 cm]

MAGNETIC PROPERTIES OF Ni - COMPLEXES IN A HYDRAZONE STRUCTURE

E. Mosiniewicz-Szablewska^{a}, A. Ślawska-Waniewska^{a}, K. Światek^{a}, M. Golebiewski^{b}

aInstitute of Physics, Polish Academy of Sciences, Al. Lotników 32, 02-668 Warsaw,

Poland bInstitute of Industrial Organic Chemistry, Annopol 6, 03-236 Warsaw, Poland

The complex of 5-methylsalicylaldehyde 5’-chlorosalicyloyl hydrazone doped with Ni^{2+} ions has been studied with electron paramagnetic resonance (EPR) spectrometer and vibrating sample magnetometer (VSM) over the temperature range 2 - 300 K. The magnetization measurements, performed in the magnetic field of 1T, show that at low temperatures the Ni^{2+} ions are coupled by the antiferromagnetic exchange interactions. Above the Néel temperature T_{N} ≅ 11.5K, the temperature dependendence of susceptibility of Ni - complexes follows the Curie - Weiss law. The EPR spectrum observed at room temperature exhibits a broad anisotropic signal originating from Ni^{2+} ions. It is also found that at low temperatures the EPR signal experiences a sharp broadening and shifts to the lower field region. The intensity of this signal decreased with lowering the temperature and below 11 K this line disappears. Instead of that, additional lines with different resonance fields can be observed. It may be attributed to the formation of an antiferromagnetically ordered phase at low temperatures.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr Ewa Mosiniewicz-Szablewska

Address for correspondence : Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46 02-668 Warsaw, Poland

Email address : mosin@ifpan.edu.pl